if person A can paint a house for 6 hours and person B can paint a house for 4 hours. How many hours they can finish in one day if they will join together painting a house?
Answer by ReadingBoosters(3246)

Rate(A) = 1/6
Rate(B) = 1/4
Rate(A+B) = 1/6 + 1/4 = 1/x
where x is the hours working together
...
Create a common denominator
4/24 + 6/24 = 10/24 = 1/x
The inverse of 1/x is x, so invert the other side of the equation also
x = 24/10 = 12/5 = 2 2/5 hours or 2.4 hours
..
Alternatively,
10/24 = 1/x
Multiply both sides by x
x(10/24) = 1
Multiply both sides by 24/10
x = 24/10 = 2.4 hours
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com